I need help I am so confused. I need to solve for k? if x=sqrt(2t/k)
Answer by stanbon(75887)

solve for k? if x=sqrt(2t/k)
-------
Square both sides to get:
(2t/k) = x^2
====
k*x^2 = 2t
-----
k = (2t)/x^2
==============
Cheers,
Stan H.
==============